Let A be a singular row reduced echelon square matrix of size n .Find a matr xA(x)=(a_(ij)(x))_(n times n) where a_(ij)(x)' s are polynomials in x such that A(0)= (a_(ij)(0))_(n times n)=A but A(delta)=(a_(ij)(delta))_(n times n) is nonsingular for all delta!=0.


1
Expert's answer
2021-06-12T07:19:00-0400

If the reduced row echelon form of a square matrix is not the identity then it can never be Non - singular

We will consider a real example below

"A=\begin{bmatrix}
 1 & 2 & -2 \\
 2 & -3 & 10\\
 2 & -3 & 10
\end{bmatrix}" This matrix can be reduced in echelon square form.


"A=\begin{bmatrix}
 1 & 2 & -2 \\
 0 & -7 & 6\\
 0 & -7 & 14
\end{bmatrix}" "R_2 \to R_2-2R_1" and "R_3 \to R_3-2R_1"


"A=\begin{bmatrix}
 1 & 2 & -2 \\
 0 & -7 & 6\\
 0 & 0 & 8
\end{bmatrix}" "R_3 \to R_3-R_2"

So here A has three pivot columns 1,-7,8. All the columns are linearly independent

"Pan \space k (A)=3 \implies |A|\not =0"

Therefore, A is non-singular.